# Edge Formation Mechanism of Single-Walled Carbon Nanotube Revealed by Molecular Dynamics Simulations based on a Neural Network Potential

### v3

@@ 2.2.1 Defect free growth of SWCNTs
Using the developed NNP, we carried out SWCNT growth simulations by supplying C atoms to Fe nanoparticles in the same way as the previous SWCNT growth simulations30, 20. We firstly performed twenty MD simulations of SWCNT growth on Fe55, Fe80, Fe100 and Fe120 nanoparticles at $T$=$1100$-$1500$ K in 100 K increments to investigate the growth trends by temperature (Figure S5). Figure 3 shows the typical growth on Fe120 nanoparticles at (a) 1100 K, and (b) 1500 K. At lower temperatures (1100-1300 K), the mobility of carbon atoms was limited, preventing their incorporation into a single large graphitic island. In these conditions, isolated hexagons nucleated simultaneously at several sites on the surface. Multiple graphitic islands then formed from these hexagons, covering the surface of the nanoparticles without lift-off and leading to the deactivation. Conversely, at high temperatures (1400-1500 K), a single hexagon was formed, enabling the sequential addition of carbon atoms to create a unified large island and the cap formation. Note that the threshold temperature is supposed to depend on the carbon supply rate, and may differ from the experimental value because the supply rate in this simulation is significantly higher than experiments.
@@
Figure 4a-h shows the SWCNT growth process at 1500 K on Fe120. At this temperature, SWCNTs grew through the typical growth stages; carbon saturation, cap formation, and side wall elongation, as reported in computational30, 20 and experimental49 studies. In the saturation stage, supplied C atoms in the gas phase were adsorbed on the catalyst surface and preferentially occupied the subsurface sites (Figure 4a). When the subsurface sites were fully occupied, carbon atoms formed small molecules such as dimer and trimer on the surface (Figure 4b). The number of small molecules such as dimers and trimers continued to increase, resulting in the carbon supersaturation of the catalyst (see Figure 4i).
In the cap formation stage, the small molecules started bonding and forming short chains as the catalyst supersaturated. The cyclization of the chain led to the first ring formation (Figure 4c). Note that as a trend over the catalyst sizes from Fe55 to Fe120, the first ring is likely to be a hexagon, which is contrary to the previous quantum mechanical molecular dynamics (QM/MD) simulations50. Additional polygons were formed next to the first hexagon and a small graphitic island was formed (Figure 4d). The island was enlarged as C atoms were supplied (Figure 4e,f), and the sixth pentagon formation finished the cap formation (Figure 4g). The number of these small molecules such as dimer and trimer continued to decrease as they transformed into rings, gradually resolving the supersaturation of the catalyst during this stage (see Figure 4i).
@@
Figure 5a, and b compare the time evolution of the numbers of hexagons, pentagons, and heptagons. The number of hexagons increased at a constant rate after the first formation. Note that in the case of the (8,7) SWCNT, the hexagon formation rate was 0.61 ring/ns and the growth rate in length was $10^{4}$ $\mu$m/s, which is about $10^{2}$ times higher than the experimental growth rates of SWCNTs51 and about 3 times higher than that of the previous SWCNT growth simulation performed using the classical interatomic potential by some of us20. The number of pentagons increased to six during the cap formation and then fluctuated but never fell below six during the sidewall elongation, indicating that the cap’s pentagons were maintained. The number of heptagons was almost always zero; even when it reached one, it immediately returned to zero. Figure 5c, and d show the detailed mechanism of defect healing on the SWCNT edge. A pentagon was formed on the edge and healed into hexagons by adding a monomer and C-C bond reformation (Figure 5c). A heptagon was formed less frequently than a pentagon, and quickly healed to a hexagon by rejecting a monomer (Figure 5d). Such defect healing close to the nanotube-catalyst interface resulted in defect-free SWCNT growth.
@@ 2.2.2 Statistical analysis of edge configurations
During the growth of CNTs, the edge configuration dynamically changed depending on the local structure such as armchair and zigzag edges. According to the thermodynamic model proposed in previous studies10, 52, the probability distribution of the edge configurations of a nanotube is governed by the interfacial Gibbs energy between the nanotube and the catalyst. To verify the consistency between the MD simulations and the thermodynamic model, we compared the probability distribution of appearance of edge configurations of a nanotube predicted by the thermodynamic model and frequency distribution obtained from the simulated SWCNT growth.
First, we classify the edge atoms in an $(n,m)$ nanotube into armchair, zigzag, antiarmchair, and antizigzag atoms, as shown in Figure 6a. If the $(n,m)$ nanotube has the smallest number of edge atoms without vacancies and adatoms, and has $i$ armchair atom pairs, the number of edge atoms in each category satisfies the following equations:
@@
The edge configurations of an $(n,m)$ nanotube are uniquely expressed by permutations of $n$ number of $a_{1}$ vectors and $m$ number of $a_{2}$ vectors. The number of unique edges of an $(n,m)$ nanotube $N(n,m)$ is thus calculated as the number of permutations in the following form:
$N(n,m)=\binom{n+m}{n}=\frac{(n+m)!}{n!m!}.$ (14)
In a previously proposed thermodynamic model52, the number of edges of an $(n,m)$ nanotube with $i$ armchair pairs $g(n,m,i)$ was calculated as the number of permutations of $i$ armchair sites and $n+m-2i$ zigzag sites:
@@
but this leads to an inequality
$N(n,m)>\sum_{i}^{m}g_{1}(n,m,i),$ (16)
because two edge configurations expressed by different permutations of vectors may be indistinguishable (see Figure S6, Supporting Information). For considering all unique edge configurations, we instead counted the number of edges using the following equation:
@@
$N(n,m)=\sum_{i}^{m}g_{2}(n,m,i).$ (18)
The derivation of $g_{2}(n,m,i)$ is explained in Supporting Information and Figure S7. Figure 6d shows the probability distribution $P(n,m,i)$ with $\Delta E_{\mathrm{Int}}=-0.131$ eV and $T=$ 300, 900, 1500, and 3000 K. The frequency distribution obtained from the simulated (8,7) SWCNT growth at 1500 K is also shown for comparison. The predicted distribution at 1500 K is consistent with the frequency distribution from the growth simulation at 1500 K in terms of the order of probability of appearance. The same analysis was performed on the (9,9) SWCNT, as shown in Figure S8. For the (9,9) SWCNT, $\Delta E_{\mathrm{Int}}$ was calculated as $-0.083$ eV. As in case of the (8,7) SWCNT, the predicted distribution $P(n,m,i)$ at 1500 K is also consistent with the frequency distribution obtained from the (9,9) SWCNT growth simulation. The temperature dependence and the interfacial energy dependence of $P(n,m,i)$ calculated with $g_{1}$ and $g_{2}$ for various chiralities are shown in Figure S9 and Figure S10, respectively. The difference in counting edges leads to qualitative differences in the temperature and interfacial energy dependence of $P(n,m,i)$. As a general trend, the diversity of edge configurations decreases with decreasing temperature, and as the interfacial energy approaches zero, the probability distribution of edge appearance approaches a distribution corresponding to the number of edges. Note that on the assumption that nanotubes grow with the edge configurations satisfying Eq.6-8, $(n,1)$ nanotubes should grow by the kink-running process regardless of the values of $T$ and $\Delta E$ as reported in the previous MD simulations by Yoshikawa et. al.20 because only edge configurations with one antiarmchair site are allowed, as shown in Figure S9 and S10.
@@ 2.2.3 Effect of ring formation at antizigzag sites
On armchair edges, pentagons were quickly healed to hexagons by adding monomers because such pentagons had dangling bonds and were highly reactive, as shown in Figure 5a. However, on zigzag edges, pentagons without dangling bonds were often formed at edge vacancies (Figure 8b). Such pentagons were less reactive, resulting in less susceptibility to monomer addition and more chances of being left unhealed inside the sidewalls. Heptagons were often formed below the pentagons to compensate for the curvature change by the pentagon formation, resulting in pairs of adjacent pentagons and heptagons (5-7 defects). 5-7 defects caused chirality changes as reported in previous experimental studies53, 54, 26. An SWCNT with 5-7 defect-mediated chirality change is shown in Figure 8a. The SWCNT grew to a certain length with the initial chirality (11,7) determined by the cap and changed to (10,7) by a 5-7 defect. We also observed the chirality change from (10,3) to (8,3) by two 5-7 defects (Figure S11).
@@ 3 Conclusions
We demonstrated defect-free SWCNT growth by MD simulation using the NNP. During the growth process, defects were frequently formed on the open edge and healed into hexagons by metal-mediated C-C bond recombination, resulting in defect-free growth. The distribution of the edge atoms during the growth was consistent with the previous thermodynamic models. Furthermore, it was shown how 5-7 defects were formed at edge vacancies, which remained unhealed in the sidewall and caused the chirality change. Also, it was observed that adatom diffusion contributed to keeping smooth edges without vacancies, which explains how a pre-determined chirality was maintained during the elongation of SWCNTs, a key mechanism for the chirality selectivity.
@@ 4 Method
To collect dataset structures covering a wide range of PESs, we sampled three types of atomic structures: disordered structures, defective graphene, and CNT growth structures. The disordered structures were obtained by randomly placing atoms in a computational cell of 8-20 Å per side, heating them to 10000 K by classical MD simulation using a Tersoff-type potential30, and then quenching them to 1000-5000 K (Figure S1a). The defective graphene structures were collected by preparing graphene in a hexagonal supercell of 17.24$\times$17.24$\times$10.0 Å, randomly displacing carbon atoms by 0-4 Å, and relaxing by MD simulations using Tersoff-type potential at 1000 K (Figure S1b). The CNT growth structures were sampled from MD simulations of SWCNT growth on Fe55 using Tersoff-type potential and NNP under development (Figure S1c). The cell size of the CNT growth structure was set to 15.0$\times$15.0$\times$15.0 Å so that the length of the vacuum space is about 10 Å. The numbers of atomic configurations included in the disordered, defective graphene and CNT growth datasets are 21744, 16871, and 5776, respectively. Each dataset was randomly separated into three subsets, with 10% as the validation set, 10% as the testing set, and the remainder as the training set.
@@
Deep potential29, one of the high dimensional NNPs, was used for training the NNP. We used the ”se_e2_a” descriptor as implemented in DeePMD-kit62. The cutoff radius is $4.0$. The smoothing function was set between 3.5-4.0 Å. For the embedding net, the sizes of the hidden layers from the input end to the output end are 25, 50, and 100, respectively. For the fitting net, three hidden layers with 240 neurons were used. Parameter optimization was performed in 2,000,000 steps. An exponentially decaying learning rate from 0.001 to 3.51×$\times 10^{-8}$ was applied. Comparisons of the energies and the forces between DFT and NNP are shown in Figure S2. The MAEs of the energies are 16.9 meV/atom and 16.1 meV/atom for the training set and testing set, and the MAEs of the forces are 0.268 eV/Å and 0.248 eV/Å for the training set and testing set, respectively.
@@
MD simulations were performed using LAMMPS63. The time step for numerical integration was 0.5 fs. The temperature of Fe nanoparticles was controlled at $T$ using the Nose-Hoover thermostat64, 65. The simulation box was a cubic cell with a size of 200$\times$200$\times$200 Å. The periodic boundary condition was applied for each direction of the axis. At the beginning of the simulation, we placed Fe55-120 nanoparticles at the center of the cell. After relaxation at constant temperature for 2 ns, we supplied C atoms at random positions in the cell controlling the density of C atoms in the gas phase at 8 atoms/cell. The initial velocity vectors of C atoms $\bm{v}$ were randomly determined, assuming that the temperature of the system is $T$ and the magnitude of the initial vector $|\bm{v}|$ satisfies $|\bm{v}|=\sqrt{3k_{b}T/m}$. In order to describe the energy barrier of the carbon feedstock decomposition, a Lennard-Jones potential was applied between non-covalent C atoms, as in previous SWCNT growth simulations30, 20. OVITO66 was used for the visualization of the atomic structures and the common neighbor analysis.
@@
I. K. conceived the project, developed the NNP, carried out the MD simulations, and analyzed the data. All authors discussed the results. I. K. wrote the original manuscript, and all authors reviewed and edited it.
@@
We thank Dr. Daniel Hedman for helpful discussion. We acknowledge Center for Computational Materials Science, Institute for Materials Research, Tohoku University for the use of MASAMUNE-IMR. A part of this work was supported by JSPS (KAKENHI JP21KK0087, JP22H01411, JP23H00174, JP23H05443) and JST (CREST JPMJCR20B5).

### v5

@@ 2.2.1 Defect free growth of SWCNTs
Using the developed NNP, we carried out SWCNT growth simulations by supplying C atoms to Fe nanoparticles in the same way as the previous SWCNT growth simulations31, 21. We firstly performed twenty MD simulations of SWCNT growth on Fe55, Fe80, Fe100 and Fe120 nanoparticles at $T$=$1100$-$1500$ K in 100 K increments to investigate the growth trends by temperature (Figure S5). Figure 3 shows the typical growth on Fe120 nanoparticles at (a) 1100 K, and (b) 1500 K. At lower temperatures (1100-1300 K), the mobility of carbon atoms was limited, preventing their incorporation into a single large graphitic island. In these conditions, isolated hexagons nucleated simultaneously at several sites on the surface. Multiple graphitic islands then formed from these hexagons, covering the surface of the nanoparticles without lift-off and leading to the deactivation. Conversely, at high temperatures (1400-1500 K), a single hexagon was formed, enabling the sequential addition of carbon atoms to create a unified large island and the cap formation. Note that the threshold temperature is supposed to depend on the carbon supply rate, and may differ from the experimental value because the supply rate in this simulation is significantly higher than experiments.
@@
Figure 4a-h shows the SWCNT growth process at 1500 K on Fe120. At this temperature, SWCNTs grew through the typical growth stages; carbon saturation, cap formation, and side wall elongation, as reported in computational31, 21 and experimental50 studies. In the saturation stage, supplied C atoms in the gas phase were adsorbed on the catalyst surface and preferentially occupied the subsurface sites (Figure 4a). When the subsurface sites were fully occupied, carbon atoms formed small molecules such as dimer and trimer on the surface (Figure 4b). The number of small molecules such as dimers and trimers continued to increase, resulting in the carbon supersaturation of the catalyst (see Figure 4i).
In the cap formation stage, the small molecules started bonding and forming short chains as the catalyst supersaturated. The cyclization of the chain led to the first ring formation (Figure 4c). Note that as a trend over the catalyst sizes from Fe55 to Fe120, the first ring is likely to be a hexagon, which is contrary to the previous quantum mechanical molecular dynamics (QM/MD) simulations51. Additional polygons were formed next to the first hexagon and a small graphitic island was formed (Figure 4d). The island was enlarged as C atoms were supplied (Figure 4e,f), and the sixth pentagon formation finished the cap formation (Figure 4g). The number of these small molecules such as dimer and trimer continued to decrease as they transformed into rings, gradually resolving the supersaturation of the catalyst during this stage (see Figure 4i).
@@
Figure 5a, and b compare the time evolution of the numbers of hexagons, pentagons, and heptagons. The number of hexagons increased at a constant rate after the first formation. Note that in the case of the (8,7) SWCNT, the hexagon formation rate was 0.61 ring/ns and the growth rate in length was $10^{4}$ $\mu$m/s, which is about $10^{2}$ times higher than the experimental growth rates of SWCNTs52 and about 3 times higher than that of the previous SWCNT growth simulation performed using the classical interatomic potential by some of us21. The number of pentagons increased to six during the cap formation and then fluctuated but never fell below six during the sidewall elongation, indicating that the cap’s pentagons were maintained. The number of heptagons was almost always zero; even when it reached one, it immediately returned to zero. Figure 5c, and d show the detailed mechanism of defect healing on the SWCNT edge. A pentagon was formed on the edge and healed into hexagons by adding a monomer and C-C bond reformation (Figure 5c). A heptagon was formed less frequently than a pentagon, and quickly healed to a hexagon by rejecting a monomer (Figure 5d). Such defect healing close to the nanotube-catalyst interface resulted in defect-free SWCNT growth.
@@ 2.2.2 Statistical analysis of edge configurations
During the growth of CNTs, the edge configuration dynamically changed depending on the local structure such as armchair and zigzag edges. According to the thermodynamic model proposed in previous studies11, 53, the probability distribution of the edge configurations of a nanotube is governed by the interfacial Gibbs energy between the nanotube and the catalyst. To verify the consistency between the MD simulations and the thermodynamic model, we compared the probability distribution of appearance of edge configurations of a nanotube predicted by the thermodynamic model and frequency distribution obtained from the simulated SWCNT growth.
First, we classify the edge atoms in an $(n,m)$ nanotube into armchair, zigzag, antiarmchair, and antizigzag atoms, as shown in Figure 6a. If the nanotube has the smallest number of edge atoms without vacancies and adatoms, $(n+m)$, and has $i$ armchair atom pairs, the number of edge atoms in each category satisfies the following equations:
@@
The degeneracy $g(n,m,i)$, i.e. the number of edges of a $(n,m)$ nanotube with $i$ armchair pairs, depends on the number of discernible species defining the nanotube edge. In the previous models, the edge was defined as a set of zigzag atoms and armchair atoms pairs, assuming either a constant number of armchair and zigzag species 11 or letting them fluctuate to account for oblique edges with respect to the tube axis 53. In the latter, $g_{1}(n,m,i)$ was calculated as the number of combinations of $i$ armchair sites and $n+m-2i$ zigzag sites:
@@
and in the former, $g_{0}(n,m)$ was the limiting case for $i=m$.
$g_{0}(n,m)=\frac{n!}{m!(n-m)!},$ (15)
These simple approximations of $g(n,m,i)$ emphasized the role of the edge configurational entropy in carbon nanotube growth, but were incomplete in the sense that the role of antiarmchair and antizigzag species that connect armchair and zigzag ones was overlooked. It is however possible to include all these four species and obtain an exact evaluation of the degeneracy $g_{2}(n,m,i)$ by defining edge configurations as sets of $n$ $a_{1}$ vectors and $m$ $a_{2}$ vectors, the unit vectors of a graphene unit cell. The derivation of $g_{2}(n,m,i)$ is explained in Supporting Information and Figures S6 and S7. Another mathematically rigorous proof was provided by F. Ducastelle 54.
@@
$\sum_{i}^{m}g_{2}(n,m,i)=\binom{n+m}{n}=\frac{(n+m)!}{n!m!}.$ (17)
Figure 6d shows the probability distribution $P(n,m,i)$ with $\Delta E_{\mathrm{Int}}=-0.131$ eV and $T=$ 300, 900, 1500, and 3000 K. The frequency distribution obtained from the simulated (8,7) SWCNT growth at 1500 K is also shown for comparison. The predicted distribution at 1500 K is consistent with the frequency distribution from the growth simulation at 1500 K in terms of the order of probability of appearance. The same analysis was performed on the (9,9) SWCNT, as shown in Figure S8. For the (9,9) SWCNT, $\Delta E_{\mathrm{Int}}$ was calculated as $-0.083$ eV. As in case of the (8,7) SWCNT, the predicted distribution $P(n,m,i)$ at 1500 K is also consistent with the frequency distribution obtained from the (9,9) SWCNT growth simulation. The temperature dependence and the interfacial energy dependence of $P(n,m,i)$ calculated with $g_{1}$ and $g_{2}$ for various chiralities are shown in Figure S9 and Figure S10, respectively. The difference in counting edges leads to qualitative differences in the temperature and interfacial energy dependence of $P(n,m,i)$. As a general trend, as the temperature decreases, the distribution of edge configurations becomes narrower and ultimately only configurations with $m$ armchair pairs survive, as expected when the configurational entropy goes to zero.
Note that on the assumption that nanotubes grow with the edge configurations satisfying Eq.6-8, $(n,1)$ nanotubes should grow by the kink-running process regardless of the values of $T$ and $\Delta E$ as reported in the previous MD simulations by Yoshikawa et. al.21 because only edge configurations with one antiarmchair site are allowed, as shown in Figures S9 and S10.
@@ 2.2.3 Effect of ring formation at antizigzag sites
On armchair edges, pentagons were quickly healed to hexagons by adding monomers because such pentagons had dangling bonds and were highly reactive, as shown in Figure 5a. However, on zigzag edges, pentagons without dangling bonds were often formed at edge vacancies (Figure 8b). Such pentagons were less reactive, resulting in less susceptibility to monomer addition and more chances of being left unhealed inside the sidewalls. Heptagons were often formed below the pentagons to compensate for the curvature change by the pentagon formation, resulting in pairs of adjacent pentagons and heptagons (5-7 defects). 5-7 defects caused chirality changes as reported in previous experimental studies55, 56, 57. An SWCNT with 5-7 defect-mediated chirality change is shown in Figure 8a. The SWCNT grew to a certain length with the initial chirality (11,7) determined by the cap and changed to (10,7) by a 5-7 defect. We also observed the chirality change from (10,3) to (8,3) by two 5-7 defects (Figure S11).
@@ 3 Conclusions
We performed MD simulations of defect-free SWCNT growth using the NNP. The MD simulations showed the structural evolution of nanotube edges with dynamic rearrangement of edge configurations. We compared the probability of appearance of the edge configurations with the thermodynamic model reported in previous work and demonstrated that the probability can be predicted from the nanotube-catalyst interfacial energy. It was also observed that the vacancy, which is caused by the ring formation at the anti-zigzag sites, leads to the formation of defects and that the adatom diffusion along the edges contributes to the healing of vacancies. These observations explain how a nanotube edge is formed and changed during the growth and provide a clue for realizing the chirality-controlled synthesis of SWCNTs.
@@ 4 Method
To collect dataset structures covering a wide range of PESs, we sampled three types of atomic structures: disordered structures, defective graphene, and structures obtained from SWCNT growth simulations. The disordered structures were obtained by randomly placing atoms in a computational cell of 8-20 Å per side, heating them to 10000 K by classical MD simulation using a Tersoff-type potential31, and then quenching them to 1000-5000 K (Figure S1a). The defective graphene structures were collected by preparing graphene in a hexagonal supercell of 17.24$\times$17.24$\times$10.0 Å, randomly displacing carbon atoms by 0-4 Å, and relaxing them using Tersoff-type potential (Figure S1b). The CNT growth structures were sampled from MD simulations of SWCNT growth on Fe55 using Tersoff-type potential and NNP under development (Figure S1c). The cell size of the CNT growth structure was set to 15.0$\times$15.0$\times$15.0 Å so that the length of the vacuum space is about 10 Å. The numbers of atomic configurations included in the disordered, defective graphene and CNT growth datasets are 21744, 16871, and 5776, respectively. Each dataset was randomly separated into three subsets, with 10% as the validation set, 10% as the testing set, and the remainder as the training set.
@@
Deep potential29 was used for training the NNP. We used the ”se_e2_a” descriptor implemented in DeePMD-kit65. The cutoff radius is set to $4.0$. The smoothing function was set between 3.5-4.0 Å. For the embedding net, the sizes of the hidden layers from the input end to the output end are 25, 50, and 100, respectively. For the fitting net, three hidden layers with 240 neurons were used. Parameter optimization was performed in 2,000,000 steps. An exponentially decaying learning rate from 0.001 to 3.51×$\times 10^{-8}$ was applied. Comparisons of the energies and the forces between DFT and NNP are shown in Figure S2. The MAEs of the energies are 16.9 meV/atom and 16.1 meV/atom for the training set and testing set, and the MAEs of the forces are 0.268 eV/Å and 0.248 eV/Å for the training set and testing set, respectively.
@@
MD simulations were performed using LAMMPS66. The time step for numerical integration was 0.5 fs. The temperature of Fe nanoparticles was controlled at $T$ using the Nose-Hoover thermostat67, 68. The simulation box was a cubic cell with a size of 200$\times$200$\times$200 Å. The periodic boundary condition was applied for each direction of the axis. At the beginning of the simulation, we placed Fe55-120 nanoparticles at the center of the cell. After relaxation at constant temperature for 2 ns, we supplied C atoms at random positions in the cell controlling the density of C atoms in the gas phase at 8 atoms/cell. The initial velocity vectors of C atoms $\bm{v}$ were randomly determined, assuming that the temperature of the system is $T$ and the magnitude of the initial vector $|\bm{v}|$ satisfies $|\bm{v}|=\sqrt{3k_{b}T/m}$. In order to describe the energy barrier of the carbon feedstock decomposition, a Lennard-Jones potential was applied between non-covalent C atoms, as in previous SWCNT growth simulations31, 21. OVITO69 was used for the visualization of the atomic structures and the common neighbor analysis.
@@
I. K. and S. M. conceived the project. I.K. developed the NNP, carried out the MD simulations. Data was analyzed and discussed by all authors. I. K. wrote the original manuscript, and all authors reviewed and edited it.
@@
We sincerely thank Dr. François Ducastelle, researcher at ONERA, France, who passed away on July 2, 2021 for his contribution to this work. He offered the mathematically rigorous derivation of the unique edge configuration in Eq. 16. We thank Dr. Daniel Hedman for helpful discussion. We acknowledge Center for Computational Materials Science, Institute for Materials Research, Tohoku University for the use of MASAMUNE-IMR. Additional computational resources were provided by Research Center for Computational Science, Okazaki, Japan (Project: 24-IMS-C245). A part of this work was supported by JSPS (KAKENHI JP21KK0087, JP22H01411, JP23H00174, JP23H05443) and JST (CREST JPMJCR20B5).